OpenStudy (kitkat1):

A system of equations is shown below: n = 3m + 7 n – 2m = 1 What is the solution, in the form (m, n), to the system of equations?

OpenStudy (kitkat1):

I will give medals!

pooja195 (pooja195):

Use substituion

OpenStudy (anonymous):

By the first equation what is n in terms of m?

pooja195 (pooja195):

\[\huge~(3m + 7)– 2m = 1\]

pooja195 (pooja195):

I put equation 1 into equation 2

pooja195 (pooja195):

Now solve that :)

OpenStudy (kitkat1):

What do I do next?

pooja195 (pooja195):

Solve that get the m value then plug the m value into equation 1

OpenStudy (kitkat1):

how can I combine the like terms?

OpenStudy (kitkat1):

I dont know what to do

pooja195 (pooja195):

3-2=?

OpenStudy (kitkat1):

1

pooja195 (pooja195):

1m+7=1 subtract 7 from both sides

OpenStudy (kitkat1):

1m= -6

pooja195 (pooja195):

good now divide :) -6/1=?

OpenStudy (kitkat1):

-6

pooja195 (pooja195):

good so we have m=-6 n=3(-6)+7

OpenStudy (kitkat1):

-18=7 ?

pooja195 (pooja195):

no -18+7-?

OpenStudy (kitkat1):

-11

pooja195 (pooja195):

good so we have m=-6 n=-11

OpenStudy (kitkat1):

but whats the answer then?

pooja195 (pooja195):

Well thats for u to figure out. We did all the work \[\huge~m=-6~~~~~~ n=-11\]

OpenStudy (kitkat1):

okay can u help me with one more?

OpenStudy (kitkat1):

A pencil at a stationery store costs $1, and a pen costs $1.50. Stella spent $21 at the store. She bought a total of 18 items. Which system of equations can be used to find the number of pencils (x) and pens (y) she bought? (5 points) 1.5x + y = 21 x = 18y 18x + y = 21 x = 1.5y x + 1.5y = 21 x + y = 18 x + 18y = 21 x = 1.5y